# Machine Learning for Two-Sample Testing under Right-Censored Data: A Simulation Study

### v1

@@ Abstract
The focus of this study is to evaluate the effectiveness of Machine Learning (ML) methods for two-sample testing with right-censored observations. To achieve this, we develop several ML-based methods with varying architectures and implement them as two-sample tests. Each method is an ensemble (stacking) that combines predictions from classical two-sample tests. This paper presents the results of training the proposed ML methods, examines their statistical power compared to classical two-sample tests, analyzes the distribution of test statistics for the proposed methods when the null hypothesis is true, and evaluates the significance of the features incorporated into the proposed methods. All results from numerical experiments were obtained from a synthetic dataset generated using the Smirnov transform (Inverse Transform Sampling) and replicated multiple times through Monte Carlo simulation. To test the two-sample problem with right-censored observations, one can use the proposed two-sample methods. All necessary materials (source code, example scripts, dataset, and samples) are available on GitHub and Hugging Face.
@@ 1 Introduction
Two-sample problem testing (and hypothesis testing in general) on censored data is an integral part of solving problems in reliability, physics, biology, medicine, economics, bioinformatics, sociology, engineering, and many other fields [1]. It is well known that there is no universal method preferred in all usage scenarios for this class of problems. Given the wide applicability, there is a constant need for improvement in the methods used. Machine learning methods, which are currently being applied in various domains, can help address this need [2].
In the construction of classical two-sample tests, reliance is primarily placed on certain parametric or non-parametric assumptions [3]. This allows the method to effectively distinguish between certain competing hypotheses, but unfortunately, it may struggle to differentiate others. Researchers have made numerous attempts to combine the properties of different tests [4, 5, 6]. However, such attempts are complicated by the peculiarities of the subject area, which involves censored data. In the literature, there are also modern works that contrapose ML methods against classical two-sample tests [7] or for the uncensored case [8, 9, 10]. However, these comparisons often involve a small number of methods. In our work, we propose considering a wide range of diverse two-sample tests. Each two-sample tests is treated as an feature, and through the use of ML-methods, we create a stacking (ensemble) that combines these features.
Therefore, the main objectives of this research are as follows:
@@
    Develop methods for two-sample problem under right-censored data using ML-methods;
@@
    Evaluate the effectiveness of the proposed methods compared to existing classical methods;
@@
    Describe the practical applicability of the proposed methods.
@@
    Publish the repositories and results of the proposed methods for public use.
@@ 2 Related Works
One group of such tests includes various generalizations of classical tests such as the Cramér–von-Mises and Kolmogorov-Smirnov for censored data, as shown in the works of Gill [11], Fleming et al. [12], Koziol [13], Schumacher [14], and Fleming [15].
Another group of tests is based on the use of weight functions, such as the weighted log-rank test. The use of different weight functions generates a multitude of different two-sample tests with new properties compared to the original ones. For example, Gehan [16], Peto and Peto [17], Tarone and Ware [18], Prentice [19], Kalbfleisch and Prentice [20], Fleming and Harrington [21, 22].
The third group of tests is based on the assumption of a specific form of competing hypotheses. This approach was used in the works of Bagdonavičius and Nikulin. They assumed that competing hypotheses could correspond to models with single (SCE-model [23]) or multiple (MCE-model [24]) intersections, and the GPH-model [25].
Additionally, due to their simplicity and efficiency, the method of constructing methods using auxiliary functions has gained popularity. These methods involve computation the value of the selection function in advance, and then using this value to make decisions about the value of the test statistic. For example, Martinez and Naranjo [5], as well as Philonenko and Postovalov [26], propose using the selection function to choose between the log-rank and Wilcoxon tests. In another work [27], authors combine the Bagdonavičius-Nikulin and weighted Kaplan-Meier tests by the selection function.
Established the potential effectiveness of this approach, the question arises of constructing a selection function that is more sensitive to the characteristics of the underlying tests. ML methods can be helpful in this regard, as they can be considered as a selection function on a set of two-sample tests, using each test as a feature in the model.
Thus, in Section 3, we present the problem formulation and describe existing statistical methods. In Section 4, we describe the proposed methods for the two-sample problem using machine learning. In Section 5, we provide the results of numerical experiments, including the training results of the proposed methods, power study of the proposed methods, study of the test statistic distribution, and feature importance results.
@@ 3.1 Problem Statement
Further, we suppose that the elements of samples are ordered: $t_{11}<\cdots<t_{1n_{1}}$ and $t_{21}<\cdots<t_{2n_{2}}$. Denote the censoring indicators $c_{ij}$ as $c_{ij}=\begin{cases}0,t_{ij}\text{~{}is failure,}\\ 1,t_{ij}\text{~{}is censored,}\\ \end{cases}$ and the censoring rate $r_{i}$ as $r_{i}=\sum_{j=1}^{n_{i}}c_{ij}/n_{i}$.
@@ 3.2.1 Log-rank test
The log-rank test is also called as Mantel–Haenszel test or Mantel–Cox test [3, 28, 29]. The test statistic is
@@ 3.2.2 Generalizations of Wilcoxon test
where $u_{i}=\begin{cases}\hat{S}_{KM}(t_{i})+\hat{S}_{KM}(t_{k_{i}}),c=0,\\ \hat{S}_{KM}(t_{k_{i}})-1,c=1,\\ \end{cases}$ $k_{i}=max\{j|j\in\{0,...,i-1\},c_{j}=0\}$, $c_{0}=0$ and $\hat{S}_{KM}(t)$ is the Kaplan-Meier estimator [30] of $S(t)$ by the pooled sample $T$.
@@ 3.2.3 Weighted tests
Weighted log-rank test. Based on the findings of [31], we describe the statistics in a unified format, covered all previously presented tests and incorporating additional ones. The test statistic can be computed as $S_{WLG}=U^{2}/\sigma$ where
@@
Let $\hat{S}_{1}(t)$ and $\hat{S}_{2}(t)$ are the Kaplan–Meier estimates [30] of the $S_{1}(t)$ and $S_{2}(t)$ consequently, and $\hat{S}(t)$ is the Kaplan–Meier estimate of the pooled sample. For every sample $X_{i}$, we construct additional sample $\overline{X}_{i}$ changing the indicator $c_{ij}$ to the opposite $\overline{c}_{ij}=1-c_{ij}$. Thus, $\hat{S}_{1}^{C}(t)$ and $\hat{S}^{C}_{2}(t)$ are the Kaplan–Meier estimates for additional samples $\overline{X}_{1}$ and $\overline{X}_{2}$.
@@ 3.3 Alternative Hypotheses
4 Proposed ML-based Methods for Two Sample Problem
We approach the two-sample problem as a binary classification task (0 if the $H_{0}$ is true, 1 if the $H_{0}$ is rejected). Each two-sample test can be considered as an independent feature. Therefore, to solve our problem, we combine the properties of various statistical methods using a stacking architecture. For the binary classification task, we consider various machine learning methods:
  1. 1.
    Logistic Regression (denote as LogReg) implemented in sklearn;
  2. 2.
    Random Forest (denote as RF) implemented in sklearn;
  3. 3.
    Gradient Boosting Machine (denote as GBM) implemented in sklearn;
  4. 4.
    XGBoost;
  5. 5.
    CatBoost [34] proposed by Yandex;
  6. 6.
    LightAutoML [35] (denote as LAMA) proposed by Sber AI Lab.
Our method takes two right-censored samples $X_{1}$ and $X_{2}$ as input. For these samples, we compute the $p$-value for the two-sample tests described earlier, since small $p$-value rejects the null hypothesis $H_{0}$. We also include information about the sample sizes and the censoring rate in order to be more sensitive for the alternative. As output, our method returns a value in the interval $[0,1]$, which serves as the test statistic for a new two-sample test. For this method, we can estimate the test power, investigate the test statistic distribution under the null hypothesis $G(S|H_{0})$, and conduct any other experiments. We detail the process of the described method in Figure 2.
Figure 2: The implementation flow chart of the proposed methods
The test statistic of each proposed method has a right-side critical area (according to the problem statement based on the binary classification task) and the null hypothesis $H_{0}$ is rejected when the test statistic is large.
For the proposed methods, we limit the number of features at the feature engineering stage, since the goal is to add information to the methods about the properties of the two-sample tests on various sample sizes and censoring rates instead of overfitting on the selected alternatives.
In training ML models, we select the best models by maximizing the Accuracy metric. Since we have a balanced dataset (50% with the $H_{0}$ is true and 50% with the $H_{0}$ is rejected) of synthetic data, maximizing Accuracy allows us to minimize the probabilities of Type I and Type II errors. The synthetic data were simulated using the Monte Carlo replications. Also, we report other metrics for reference: the Precision, Recall, ROC AUC, Average Precision, Specificity and Proportion of predictions without FN (false negative): $(TP+FP+TN)/(TP+FP+FN+TN)$.
You also can solve the two-sample problem by the proposed methods which are available in GitHub.
@@ 5.1 Dataset
      if $X_{1}\sim\xi_{1}$ compares to $X_{2}\sim\xi_{1}$, then $H_{0}$ is true. It is indicated as ’H0’;
@@
      if $X_{1}\sim\xi_{1}$ compares to $X_{2}\sim\xi_{2}$, then $H_{0}$ is rejected. It is indicated as ’H1’;
@@ 5.3 Study of the Test Power
For values of the average rank, we also compute the values of the Wald [36] and Savage [37] criteria, which are used for decision-making under risk and uncertainty of environment. In this case, the ”environment” is an alternative hypothesis, and the utility function is the rank of the method on the alternative hypothesis. The lower the rank of the method, the higher it is in the ordered list, which means that this method is better than others. The Wald criterion focuses on maximizing utility in worst-case scenarios, whereas the Savage criterion aims to minimize ’regret’ associated with the utility loss compared to the best possible outcome. A similar methodology has already applied previously in [38].
@@ 6 Discussion
The behavior of the test statistic distribution of the proposed methods under the $H_{0}$ is true is also investigated. For this purpose, we constructed $G(S|H_{0})$ under various parameters $(n_{i},r_{i},F_{i},F^{C}_{i}),i=1,2$ and computed the average error between the middle line and the minorant/majorant of this set (in the region $0<\alpha\leq 0.1$). This shows what order of error can be expected when calculating the $p$-value. The obtained results show that Monte Carlo simulating the test statistic distribution to determine the critical value can be an effective way to solve this issue. A detailed table of percentage points was prepared.
The advantage of the proposed methods is that it is based on many classical and well-studied two-sample tests. This allows the method to be sensitive to the properties of these tests and aggregate their synergistic effect. Similar approaches based on the selection function have been used previously [4, 26, 27]. The choice of such a function was associated with taking into account the special statistical properties of the methods, which was a limitation for a wide study in this direction. ML methods allow one to be more sensitive to the properties of the underlying methods and at the same time play the role of the selection function. Note also that the obtained test power results confirm the results in [27] when studying the MIN3 test.
To apply these methods in practice, the full source code for the implementation of the proposed methods is published in the GitHub repository (including trained models and an example to run). If even this is not enough, the same repository contains the full source code for modeling the original dataset.
@@
The study the limit test statistic distribution of the proposed methods can significantly increase the attractiveness and accessibility of these methods for solving practical problems. In our opinion, this can be an one of the prioritiest task for further possible research.

### v2

@@ Abstract
The focus of this study is to evaluate the effectiveness of Machine Learning (ML) methods for two-sample testing with right-censored observations. To achieve this, we develop several ML-based methods with varying architectures and implement them as two-sample tests. Each method is an ensemble (stacking) that combines predictions from classical two-sample tests. This paper presents the results of training the proposed ML methods, examines their statistical power compared to classical two-sample tests, analyzes the null distribution of the proposed methods when the null hypothesis is true, and evaluates the significance of the features incorporated into the proposed methods. In total, this work covers 18 methods for two-sample testing under right-censored observations, including the proposed methods and classical well-studied two-sample tests. All results from numerical experiments were obtained from a synthetic dataset generated using the inverse transform sampling method and replicated multiple times through Monte Carlo simulation. To test the two-sample problem with right-censored observations, one can use the proposed two-sample methods (scripts, dataset, and models are available on GitHub and Hugging Face).
@@ 1 Introduction
Two-sample testing (and hypothesis testing in general) on censored data is integral to addressing problems across various fields, including reliability, physics, biology, medicine, economics, bioinformatics, sociology, and engineering [1]. Recent advanced studies further corroborate this, demonstrating applications of two-sample testing in exploring algorithms in engineering [2], chemical reactions [3], new medicines [4], ecosystems [5], and the reliability of engineering systems [6] and dynamic systems [7]. In other words, two-sample testing is a universal multi-domain approach applicable to a wide array of contemporary scientific challenges.
It is well known that no universally preferred method exists for this problem across all scenarios. Given its wide applicability, there is a constant challenge to improve the tests used. Classical two-sample tests primarily rely on certain parametric or non-parametric assumptions [8], enabling effective differentiation between some competing hypotheses (though they may struggle with others). Consequently, researchers have made numerous attempts to combine and study the properties of different tests [9, 10, 11]. However, such attempts are complicated by the peculiarities of the subject area, particularly concerning censored (incomplete) observations. Notably, recent works continue to develop robust tests using increasingly complex statistical assumptions and methodologies [12]. Machine learning methods, currently applied across various domains [13], can also help address this challenge. While there are modern works based on machine learning methods [14], studies often focus more on uncensored observations [15], consider a limited number of methods, and do not compare their effectiveness with classical two-sample tests [16]. Additionally, these works frequently pay insufficient attention to the statistical properties of the methods, often limiting their analysis to statistical power [17]. It is essential to note that the null distribution can significantly differ under alternatives not considered in prior works, potentially leading to changes in the binary threshold and resulting in incorrect statistical conclusions.
In summary, this is a promising new direction that has not been sufficiently explored from various angles. Therefore, in our work, we consider a wide range of classical two-sample tests applicable to right-censored data that are well-studied in the literature, widely used in practice, and possess distribution-free properties. We explore methods for the two-sample problem using these classical tests as features and forming an ensemble (stacking) for machine learning. The proposed methods are presented and investigated as standard statistical techniques. This approach allows us to combine the advantages of classical tests. For the proposed methods, we investigate power and null distribution under various alternatives using Monte Carlo simulations.
Thus, the main contributions of this study are as follows:
@@
    Develop statistical methods for two-sample testing with right-censored observations using machine learning techniques;
@@
    Evaluate the power of the proposed methods in comparison to classical tests;
@@
    Study the null distribution numerically under various alternatives to describe the practical applicability of the proposed methods.
@@
    Publish repositories of the proposed methods for public use.
@@ 2 Related Works
One group of such tests includes various generalizations of classical tests such as the Cramér–von Mises and Kolmogorov–Smirnov tests for censored data, as shown in the works of Fleming et al. [18], Koziol [19], and Schumacher [20].
Another group of tests is based on the use of weight functions, such as the weighted log-rank test. The use of different weight functions generates a multitude of different two-sample tests with new properties compared to the original ones. For example, see Gehan [21], Peto and Peto [22], Tarone and Ware [23], Prentice [24], Kalbfleisch and Prentice [25], and Fleming and Harrington [26].
The third group of tests is based on the assumption of a specific form of competing hypotheses. This approach was used in the works of Bagdonavičius and Nikulin, who assumed that competing hypotheses could correspond to models with single (SCE-model [27]) or multiple (MCE-model [28]) intersections, as well as the GPH-model [29].
Additionally, due to their simplicity and efficiency, methods that construct auxiliary functions have gained popularity. These methods involve computing the value of the selection function in advance and then using this value to make decisions about the value of the test statistic. For example, Martinez and Naranjo [9], as well as Philonenko and Postovalov [30], propose using the selection function to choose between the log-rank and Wilcoxon tests. In another work [31], the authors combine the Bagdonavičius-Nikulin and weighted Kaplan-Meier tests using the selection function.
Having established the potential effectiveness of this approach, the question arises of constructing a selection function that is more sensitive to the characteristics of the underlying tests. ML-based methods can be helpful in this regard, as they can be considered a selection function over a set of two-sample tests, using each test as a feature in the model.
Thus, in Section 3, we present the problem formulation, describe classical two-sample tests, and outline the proposed methods for the two-sample problem based on machine learning. In Section 4, we provide the results of numerical experiments, including training results for the proposed methods, a power study comparing the proposed methods to classical tests, an examination of the null distribution, and feature importance results.
@@ 3.1 Problem Statement
Further, we suppose that the elements of samples are ordered: $t_{11}<\cdots<t_{1n_{1}}$ and $t_{21}<\cdots<t_{2n_{2}}$. Denote the censoring indicators $c_{ij}$ as $c_{ij}~{}=~{}\begin{cases}0,t_{ij}\text{~{}is failure,}\\ 1,t_{ij}\text{~{}is censored,}\\ \end{cases}$ and the censoring rate $r_{i}$ as $r_{i}=\sum_{j=1}^{n_{i}}c_{ij}/n_{i}$.
@@ 3.2.1 Log-rank test
The log-rank test is also called as Mantel–Haenszel test or Mantel–Cox test. The test statistic [33, 32, 8] is
@@ 3.2.2 Generalizations of Wilcoxon test
where $u_{i}=\begin{cases}\hat{S}_{KM}(t_{i})+\hat{S}_{KM}(t_{k_{i}}),c=0,\\ \hat{S}_{KM}(t_{k_{i}})-1,c=1,\\ \end{cases}$ $k_{i}=max\{j|j\in\{0,...,i-1\},c_{j}=0\}$, $c_{0}=0$ and $\hat{S}_{KM}(t)$ is the Kaplan-Meier [34] estimator of $S(t)$ by the pooled sample $T$.
@@ 3.2.3 Weighted tests
Weighted log-rank test. Based on the findings in [27], we describe the statistics in a unified format, covered all previously presented tests and incorporating additional ones. The test statistic can be computed as $S_{WLG}=U^{2}/\sigma$ where
@@
Let $\hat{S}_{1}(t)$ and $\hat{S}_{2}(t)$ are the Kaplan–Meier estimates of the $S_{1}(t)$ and $S_{2}(t)$ consequently, and $\hat{S}(t)$ is the Kaplan–Meier estimate of the pooled sample. For every sample $X_{i}$, we construct additional sample $\overline{X}_{i}$ changing the indicator $c_{ij}$ to the opposite $\overline{c}_{ij}=1-c_{ij}$. Thus, $\hat{S}_{1}^{C}(t)$ and $\hat{S}^{C}_{2}(t)$ are the Kaplan–Meier estimates for additional samples $\overline{X}_{1}$ and $\overline{X}_{2}$.
@@ 3.2.5 Two-Stage tests
3.3 Proposed ML-based Methods for Two Sample Problem
The two-sample tests described earlier are classical methods for addressing this problem. These tests offer several advantages: (1) They are well-studied, with extensive theoretical investigations available. (2) They have been used for many years to solve applied problems (e.g., the log-rank test has been utilized for decades in medical research). (3) Almost all tests possess the important property of being distribution-free, ensuring asymptotic convergence to the corresponding limit distribution. (4) Each test is preferable under specific alternative hypotheses, where other tests may be unsuitable.
This suggests that combining these statistics can yield a test statistic that is more comprehensive and sensitive to a broader range of alternatives than each individual test statistic. Furthermore, such a combined test would be more robust and powerful. To achieve this, we consider these tests as weak classifiers that can be integrated into a strong classifier.
We formulate the two-sample problem as a binary classification task: 0 if $H_{0}$ is accepted, 1 if $H_{0}$ is rejected. Then, the classifier’s input consists of two right-censored samples, while the output is a predicted logit in [0, 1], representing the ”probability” of rejecting $H_{0}$. Each two-sample test is treated as an independent feature, allowing us to combine the properties of various statistical methods using a stacking architecture to solve the problem. Thus, a statistical conclusion $D$ through the proposed methods can be formulated as follows:
$D=\begin{cases}H_{0}~{}\text{is accepted, if}~{}p_{v}\geq\alpha\\ H_{0}~{}\text{is rejected, if}~{}p_{v}<\alpha\end{cases}$
where $p_{v}=1-G_{n_{1},n_{2},r_{1},r_{2}}(pred_{n_{1},n_{2},r_{1},r_{2}}|H_{0})$, $pred_{n_{1},n_{2},r_{1},r_{2}}=\text{Classifier}(X_{1},X_{2},\Sigma)$, $G_{n_{1},n_{2},r_{1},r_{2}}(\cdot|H_{0})$ is the pre-limit null distribution, $\alpha$ is the probability of a Type I error, $\Sigma=\Sigma_{1}\cup\Sigma_{2}$, $\Sigma_{1}=\{S_{P},S_{G},S_{LG},S_{\text{SCE}},S_{\text{MCE}},S_{\text{GPH}},S_{\text{TW}},S_{\text{PPr}},S_{\text{Pr}},S_{WKM}\}$, and $\Sigma_{2}=\{S_{Q},S_{MAX},S_{MIN3}\}$. We also detailly demonstrate the process of the described method in Figure 1. Hence, the test statistic for each proposed method $S_{\text{ML}}$ is given by $S_{\text{ML}}=pred_{n_{1},n_{2},r_{1},r_{2}}$. This representation allows for the investigation of all standard statistical procedures (e.g., test power and null distribution).
A binary Classifier$(X_{1},X_{2},\Sigma)$ needs to be trained on the target $y\in\{0,1\}$ ($y=0$ if the $H_{0}$ is satisfied for $X_{1},X_{2}$; otherwise, $y=1$) and on features that are formed as follows:
$\text{Classifier}(X_{1},X_{2},\Sigma):=\begin{cases}p\text{-value}(X_{1},X_{2},S_{l}),&\text{if }S_{l}\in\Sigma_{1},\\ S_{l}(X_{1},X_{2}),&\text{if }S_{l}\in\Sigma_{2},\\ I(\min(n_{1},n_{2})\in[n_{L},n_{R}]),\\ I(\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}\in[r_{L},r_{R}]),\end{cases}$
where $I(A)=1$ if $A$ is True; otherwise, $I(A)=0$. We set $[n_{L},n_{R}]$ to represent small 50-, low (50, 100], medium (100, 500], and high 500+ sample sizes. For $[r_{L},r_{R}]$, we categorize it as null [0%, 0%], low (0%, 15%], medium (15%, 35%], and high 35%+ censoring rates. For the proposed methods, we limit the number of features during the feature engineering stage since the goal is to enhance information about the properties of the two-sample tests rather than overfitting on the selected alternatives.
To solve the binary classification task, we apply several machine learning methods: (1) Logistic Regression (denoted as LogReg) implemented in sklearn; (2) Random Forest (denoted as RF) implemented in sklearn; (3) Gradient Boosting Machine (denoted as GBM) implemented in sklearn; (4) XGBoost; (5) CatBoost proposed by Yandex [37]; (6) LightAutoML (LAMA) proposed by Sber AI Lab [38].
For the proposed methods, the limit distribution $G(x|H_{0})$ is unknown in analytical form. However, this is not necessary for applying the statistical method, as it is sufficient to study how the pre-limit distribution $G_{n_{1},n_{2},r_{1},r_{2}}(x|H_{0})$ can significantly vary. Therefore, we numerically study the pre-limit distribution of the proposed methods and determine the possibility of their applicability in practice. To strengthen the conviction of the numerical results obtained, we use competing hypotheses of diverse distribution families $F(t)$ and $F^{C}(t)$, as described in Section 3.4. We also note that the proposed methods have a right-tailed critical area (according to the problem statement based on binary classification) and the null hypothesis $H_{0}$ is rejected when the test statistic is large.
Figure 1: Implementation flow chart of the proposed methods
In training ML models, we select the best models by maximizing the Accuracy metric. Since we have a balanced dataset (50% where $H_{0}$ is true and 50% where $H_{0}$ is rejected) of synthetic data, maximizing Accuracy allows us to minimize the probabilities of Type I and Type II errors. The synthetic data were simulated using Monte Carlo replications. Additionally, we report other metrics for reference: Precision, Recall, ROC AUC, Average Precision, Specificity, and Proportion of predictions without FN (false negative): $(TP+FP+TN)/(TP+FP+FN+TN)$.
The proposed methods can also be used to solve the two-sample problem and are available on GitHub.
@@ 4.1 Dataset
      if $X_{1}\sim\xi_{1}$ compares to $X_{2}\sim\xi_{1}$, then the $H_{0}$ is true (indicated as ’H0’);
@@
      if $X_{1}\sim\xi_{1}$ compares to $X_{2}\sim\xi_{2}$, then the $H_{0}$ is rejected (indicated as ’H1’);
@@ 4.3 Study of the Test Power
For values of the average rank, we also compute the scores of the Wald [39] and Savage [40] criteria, which are used for decision-making under risk and uncertainty of environment. In this case, the ”environment” is an alternative hypothesis, and the utility function is the rank of the method on the alternative hypothesis. The lower the rank of the method, the higher it is in the ordered list, which means that this method is better than others. The Wald criterion focuses on maximizing utility in worst-case scenarios, whereas the Savage criterion aims to minimize ’regret’ associated with the utility loss compared to the best possible outcome. The same methodology has already applied previously in [31].
@@ 5 Discussion
The behavior of the test statistic distribution of the proposed methods under the $H_{0}$ is true is also investigated. For this purpose, we simulated $G(S|H_{0})$ under various parameters $(n_{i},r_{i},F_{i},F^{C}_{i}),i=1,2$ and computed the average error between the middle line and the minorant/majorant of this set (in the region $0<\alpha\leq 0.1$). This shows what order of error can be expected when calculating the $p$-value. The obtained results show that Monte Carlo simulating the test statistic distribution to determine the critical value can be an effective way to solve this issue. A detailed table of percentage points was prepared.
The advantage of the proposed methods is that it is based on many classical and well-studied two-sample tests. This allows the method to be sensitive to the properties of these tests and aggregate their synergistic effect. Similar approaches based on the selection function have been used previously in [9, 30, 31]. The choice of such a function was associated with taking into account the special statistical properties of the methods, which was a limitation for a wide study in this direction. ML methods allow one to be more sensitive to the properties of the underlying methods and at the same time play the role of the selection function. Note also that the obtained test power results confirm the results of when studying the MIN3 test [31].
To apply these methods in practice, the full source code for the implementation of the proposed methods is published in the GitHub repository (including trained models and scripts to run). If even this is not enough, the same repository contains the full source code for modeling the original dataset.
@@
The study the limit null distribution of the proposed methods can significantly increase the attractiveness and accessibility of these methods for solving practical problems. In our opinion, this can be an one of the prioritiest task for further possible research.